Suppose 2 firms compete in a market for widgets. Each of them produces identical widgets. Each firm incurs a cost of $10 per widget produced. The two firms simultaneously (and independently) decide how many widgets to produce. The inverse demand for widgets is given by P=100−3Q, where Q=q1+q2, where q1denotes the output of firm 1 and q2 denotes the output of firm 2.

What is firm 1's best response to q2=2?

What is firm 1's best response to q2=20?

What is firm 2's output in the Nash equilibrium?
Answer #1

In this case, firms' decisions are not independent of their rivals' actions. This is a case of a Cournot simultaneous move game where firms face a common aggregate demand curve in the market, and they simultaneously select their outputs, given their marginal cost functions, as the best response corresponding to the opponents' strategy.

Here, we have the given inverse demand function in the market as P = 100 - 3Q. We can rewrite Q = q1 + q2, where q1 represents firm 1's output, and q2 represents firm 2's output.

Hence, P = 100 - 3(q1 + q2)

Since, the firms are perfectly symmetric to each other, with the same MC function, we need only find one firm's optimal output. Hence, for firm 1's optimal output:

P = (100 - 3q2) - 3q1

Remember, a firm's total revenue, which we denote by R, is calculated by multiplying price with the output produced. Therefore,

R = P.q1 = (100 - 3q2)q1 - 3q12

or R = 100q1 - 3q2q1 - 3q12 --------------- [1]

We can calculate the marginal revenue of the firm by differentiating [1] with respect to q1

MR = 100 - 3q2 - 6q1

In equilibrium, we have Marginal cost (given as $10) equated with Marginal Revenue, and thus, we get

100 - 3q2 - 6q1 = 10

or 90 - 6q1 - 3q2 = 0

or 30 - 2q1 - q2 = 0

Therefore, q1 = 15 - 0.5q2 is the best response function of firm 1 in response to firm 2 in the market. Since we noted earlier that firm 1 and firm 2 are perfectly symmetric firms, we can easily infer the following:

q2 = 15 - 0.5q1 ----------- [2]

(this can be verified easily by calculating in a similar fashion the MR of firm 2, and equating with MC =10)

(a) Firm-1's best response to q2 = 2 is q1 = 15 - 0.5(2) = 15 -1 = 14 units

(b) Firm-1's best response to q2 = 20 is q1 = 15 - 0.5(20) = 15 - 10 = 5 units

(c) To calculate Nash equilibrium, we must first understand what a Nash equilibrium is. A Nash equilibrium is an n-tuple strategy profile (a1*,......, an*) where ai* for all i = 1,..., n is the best response strategy. That is, in a Nash equilibrium, all players are playing their best response strategies.

To compute the Nash equilibrium outcome, we substitute the value of q1 in [2] and compute the value of q2 in the best response function, which is the Nash equilibrium output:

q2 = 15 - 0.5(15 - 0.5q2)

q2 = 15 - 7.5 + 0.25q2

0.75q2 = 7.5

q2 = 10 = q1 in Nash equilibrium